OpenStudy (anonymous):

Find the volume and surface area a sphere with the great circle having an area of 26pi

OpenStudy (anonymous):

great circle of a sphere = largest circle made at the diameter of the sphere. Radius of great circle = radius of sphere

OpenStudy (anonymous):

Area of circle = \(\pi r^2\) use this to get the area of sphere Volume of sphere = \({4\over 3}\pi r^3\) Surface area of sphere = \(4\pi r^2\)

OpenStudy (anonymous):

idk what is the radius

OpenStudy (anonymous):

Area of circle = 26 pi plug into the equation \[26\pi=\pi r^2\]solve for r

OpenStudy (anonymous):

r=5.1

OpenStudy (anonymous):

this is also the radius of the sphere. plug this radius in the formulae for the volume and surface area

OpenStudy (anonymous):

thank u :)